OpenStudy (anonymous):

6 lb of peanuts which cost $4/lb were combined with 3 lb of spices which cost $1/lb. find the cost per lb of the mixture.

OpenStudy (amistre64):

6(4) + 3(1) = 9lbs(x) 24 + 3 = 9lbs(x) 27 dollars = 9lbs(x) 27/9 = x 3 dollars per lb.....

OpenStudy (anonymous):

thank u